OpenStudy (anonymous):

Isolate the variable: 4+n=18+2n

OpenStudy (anonymous):

n=-14

OpenStudy (anonymous):

are you very sure

OpenStudy (anonymous):

@estherlouise Welcome to OpenStudy, but make to guide the user to the answer instead of giving out the direct answer :) It teaches the asker how to do a similar problem in the future

OpenStudy (anonymous):

so can u guys walk me through it?

TheSmartOne (thesmartone):

\(\bf\huge\color{#ff0000}{W}\color{#ff2000}{e}\color{#ff4000}{l}\color{#ff5f00}{c}\color{#ff7f00}{o}\color{#ffaa00}{m}\color{#ffd400}{e}~\color{#bfff00}{t}\color{#80ff00}{o}~\color{#00ff00}{O}\color{#00ff40}{p}\color{#00ff80}{e}\color{#00ffbf}{n}\color{#00ffff}{S}\color{#00aaff}{t}\color{#0055ff}{u}\color{#0000ff}{d}\color{#2300ff}{y}\color{#4600ff}{!}\color{#6800ff}{!}\color{#8b00ff}{!}\\\small\cal Made~by~@TheSmartOne\) Hey there!!! Since you are new here, read this legendary tutorial for new OpenStudiers!! http://openstudy.com/study#/updates/543de42fe4b0b3c6e146b5e8 @MSxSammiBooxX will be helping you! Just thought to welcome you and tell you not to give direct answers. Because you will get suspended but since you are new its ok. Just don't do it next time!

OpenStudy (anonymous):

Yes :) Let's re-arrange the problem to get this, it'll make things easier n + 4 = 2n + 18 Anyways we can start this by subtracting 2n on both sides of the equation, like so n + 4 - 2n = 2n + 18 - 2n Now we have -n + 4 = 18... Do you know how to solve this?

OpenStudy (anonymous):

yes you isolate the variable witch means you subtract 4 from both sides -n=14

OpenStudy (anonymous):

so would that be the answer?

OpenStudy (anonymous):

Correct and yes :)

OpenStudy (anonymous):

ok thanks so much! medal to you :D

OpenStudy (anonymous):

No problem, and thanks! If you ever need more help tag me @MSxSammiBooxX or @sammixboo when that account is online :)

OpenStudy (anonymous):

ok thanks